OpenStudy (anonymous):

A boy finds an abandoned mine shaft in the woods, and wants to know how deep the hole is. He drops in a stone, and counts 4 seconds before he hears the "plunk" of the stone hitting the bottom of the shaft. Approximately how deep is the shaft? Assume a gravitational acceleration of 9.8 m/s2 and no significant air resistance hint:.(Hint: d=at2/2)

OpenStudy (shamim):

h=ut+(1/2)at^2=0*t+(1/2)gt^2=0.5*9.8*t^2=0.5*9.8*4^2

OpenStudy (shamim):

tell me h=?

OpenStudy (anonymous):

14 meters 24 meters 39 meters 78 meters its between these answers

OpenStudy (shamim):

ya answer is in ur optoins

OpenStudy (shamim):

plz use ur calculator to get the right answer

OpenStudy (anonymous):

tell me how to start

OpenStudy (shamim):

0.5*9.8*16=?

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

did u get the result?

OpenStudy (shamim):

thanks for giving medal

OpenStudy (anonymous):

no problemo